OpenStudy (anonymous):

I have points (-2, -1) (1, 2) and I need to write an equation in point-slope form, one in standard form, and one in slope-intercept form...help please?

OpenStudy (binary3i):

Y=mX+c where m is the slope and c is the Y intersept. m=(Y1-Y2)/(X1-X2)=(-2-1)/(-1-2)=1 therefore m=1 put this in to equation. Y=X+c using one of the point put Y=2 and X=1 and get the c 2=1+c c=1 now your done and you equation is Y=X+1

OpenStudy (anonymous):

which one is that?

OpenStudy (binary3i):

Y=X+1